How do you solve (x+b)(x-b)=x(x-5) for x? I would like to see every step and it would be nice if you could explain why you did each step. Thank you very much.
Answer by stanbon(57312)

solve (x+b)(x-b)=x(x-5) for x
Use the distributive law to get;
(x+b)x - (x+b)b = x^2-5x
Use the distributive law again to get:
x^2+bx-bx-b^2=x^2-5x
Simplify the left side to get:
x^2-b^2=x^2-5x
Subtract x^2 from both sides to get:
-b^2=-5x
Divide both sides by -5 to get:
x=b^2/5
Cheers,
Stan H.